OpenStudy (rmrjr22):

If a basketball player, with 3 seconds remaining in the game, throws the ball at a 45 degree angle from 4 feet off the ground, standing 70 feet from the basket, which is 10 feet in the air, a) what initial velocity does the ball need to have in order to reach the basket? b) What is the time of flight? c) How much time, if any, will be left in the game after the shot?

OpenStudy (anonymous):

 Always draw first :) 1) Find t in term of v0 with this eq: \[x=v_0\cos{45}*t\] 2) Plud this t into \[y=y_0+v_0\sin{45}*t-\frac12 gt^2\] To find v0. That's answer for a. As for b: (Just do a first)

OpenStudy (rmrjr22):

I ended up with \[V _{0} = 26.38 ft/s\]

OpenStudy (anonymous):

I haven't tried it yet. If you did all correctly, including unit sync, you'll come to correct answer. Can you find b) now??

OpenStudy (rmrjr22):

\[70 = 0 + V_{0} \cos \Theta t \]

OpenStudy (rmrjr22):

\[t= 70/(.707(26.38)) = 3.75 s\]

OpenStudy (anonymous):

Correct :)

OpenStudy (rmrjr22):

Only problem is my professor ended up getting 50.16 ft/s then 1.9739 s

OpenStudy (anonymous):

what g did you use?

OpenStudy (rmrjr22):

i used 9.8m/s^{2}... he used 32.2ft?^2

OpenStudy (rmrjr22):

32.2ft/s^2***

OpenStudy (anonymous):

As I said, you must sync the unit. Here, for consistency, we use g=32.2 ft/s^2. Because all other length unit are in ft.

OpenStudy (rmrjr22):

i dont understand "sync the unit"?

OpenStudy (anonymous):

It means consistency in unit. I don't know better term :)

OpenStudy (rmrjr22):

haha nvr mind. Got it as soon as I was typing it out. Its 9.8 m/s^s = 32ft/s^2

OpenStudy (anonymous):

Fine then :)